OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±8) and asymptotes at y equals inverse of plus minus 1 divided by 2x

OpenStudy (ybarrap):

Using sketch below: $$\bf \dfrac{y^2}{a^2}-\dfrac{x^2}{b^2}=1 =(^\star)\\ a=\pm 8\\ \frac{c}{a}=\cos x \implies c=a\cos x\\ b=\sqrt{c^2-a^2}\\ \text{Plug-in values for a,b and c into }(^\star) $$